# Neutron skins probed in proton knockout from neutron-rich nuclei

### v1

@@ I Introduction
Fragmentation and knockout reactions, such as (p,2p) (and possibly (p,3p)), at intermediate and high energies could provide a potentially sensitive probe of neutron skins Kox1987 ; Zenihiro2018 . In these reactions, initial- and final-state interactions depend mainly on the total density $\rho=\rho_{p}+\rho_{n}$ and are strongly surface weighted due to absorption and attenuation. Sensitivity to neutron densities is enhanced because $\sigma_{pn}>\sigma_{pp}$ at several hundred MeV to GeV energies, making reaction dynamics increasingly neutron dominated. Glauber-type attenuation, nuclear transparency, and multiple-scattering effects therefore depend primarily on the surface matter distribution rather than on volume-averaged densities. This neutron dominance is most pronounced in the intermediate-energy regime where Glauber and eikonal descriptions are applicable and tends to saturate at asymptotically higher energies. Consequently, incorrect neutron densities can bias extracted spectroscopic factors, nuclear radii, and reaction cross sections by 5–20% or more, primarily through reaction filtering and absorption rather than through bound-state normalization. Glauber and DWIA studies have shown that reaction observables such as $(p,2p)$ cross sections, nuclear transparency, and knockout spectroscopic factors require accurate neutron densities to be interpreted reliably Hufner1985 ; Giusti1999 ; Tostevin1999 ; Hen2017 ; Ogata2015 .
Momentum distributions are particularly sensitive to the spatial extent of nucleon wave functions, and therefore can reflect surface properties such as neutron skins and halos. In neutron-rich nuclei, valence neutrons may occupy weakly bound orbitals extending into the low-density surface region. The increased spatial localization length leads, via the uncertainty principle, to narrower momentum distributions. As a result, knockout momentum distributions provide a clean probe of density tails and surface diffuseness, complementing information obtained from interaction cross sections or elastic scattering. At higher missing momenta, typically above $p\sim 250$$300\penalty 10000\ \mathrm{MeV}/c$, momentum distributions probe physics beyond the independent-particle picture. In this regime, contributions from short-range and tensor correlations dominate, giving rise to high-momentum tails that are absent in mean-field models. Measurements of these tails, and their isospin dependence in $(p,2p)$ versus $(p,pn)$ reactions, provide direct evidence for correlated nucleon pairs and help explain the systematic quenching of spectroscopic factors observed in knockout experiments.
@@
Ref. Aum13 has shown that experimentally observed momentum distributions are intrinsically reaction-dependent quantities, reflecting a nontrivial convolution of nuclear structure and reaction dynamics rather than a simple Fourier transform of the bound-state wave function. Reaction distortions further modify the shape of momentum distributions. Initial- and final-state interactions, recoil effects, and off-shell kinematics can lead to broadening, skewness, or asymmetries, even for nominally symmetric orbitals Giusti1999 ; Aum13 ; Ryckebusch2003 . An additional complication arises in the interpretation of high-momentum components. While short-range and tensor correlations in the nuclear wave function generate genuine high-momentum tails, similar features can also emerge from multistep scattering and rescattering processes inherent to the reaction mechanism. Without a consistent reaction model, it is therefore not possible to attribute high-momentum strength uniquely to many-body correlations. This ambiguity is particularly relevant for $(p,2p)$ and $(p,pn)$ reactions at GeV energies, where both correlated dynamics and reaction-induced effects may contribute Hen2017 ; subedi:2008:SCIENCE .
@@ II Cross section formalism
Another problem is that the “straight-line approximation”, implicit in the Glauber method, Eqs. (3-6), is not adequate in the treatment of specific process and their details, such as in (p,2p) and (p,3p) reactions. The physical picture of the most commonly adopted Glauber formalism for proton-nucleus collisions is that the incoming proton travels straight through the nucleus at high energy and at each transverse position $\mathbf{b}$, the probability of interacting with nucleons is described by the product of survival factors $\left[1-\Gamma_{pN}\right]$. Multiple scattering effects arise from the product over all nucleons: the proton can scatter off one, two, or more nucleons during its passage along a straight-line. Although this approximation neglects deflection and quantum interference effects, it has been shown to provide a reasonable leading-order description of inclusive observables at intermediate and high energies ($\gtrsim 200$ MeV/nucleon).
@@
between two binary collisions happening between positions ${\bf r}_{1}=({\bf b},z_{1})$ and ${\bf r}_{2}=({\bf b},z_{2})$. Between the two positions, the leading particle is assumed to move along a straight-line path with transverse coordinate $\bf b$}. After one collision, the protons are allowed to move in a different direction. As long as their energies are large enough, they can be appropriately described by eikonal wavefunctions. Their wavefunctions may be further influenced by a real mean-field potential $U_{R}$. However, instead of amplitudes, as done in Ref. Aum13 , we will work with the probabilistic reaction propagator
@@
The wavefunctions squared, $|\psi_{\alpha}({\bf r})|^{2}$, account for the knockout of a proton from an orbital with quantum numbers defined by $\alpha$. The functions $W(\Omega)$ are the likelihood for scattering at angles $\Omega$ in the c.m. of the colliding pair. In the case of elastic collisions,
@@
with $d\sigma_{el}/d\Omega$ and $\sigma_{el}$ being the differential and total elastic scattering cross sections, without Coulomb interaction. In summary (see Fig. 1), The impact parameter ${\bf b}_{1}$, and ${\bf b}_{1}^{\prime}$, corresponds to the vector perpendicular distance between the nuclear center and the velocity vector for the leading proton, and of the proton knocked out from the nucleus, after the first collision. A similar notation, with ${\bf b}_{2}$ and ${\bf b}_{2}^{\prime}$, is used for the second collision. The spectroscopic information about the removed protons are encoded in the wavefunctions $\psi_{i}({\bf r}_{j})$. This formalism could be easily extended to inelastic collisions such as pion production, by replacing in $W(\Omega)$ the cross sections for pion production, or any other particle produced during the collisions. For (p,2p) collisions, we will use a similar method, with
@@
One can also extend it to more than two binary collisions. In contrast to the method employed in Ref. Aumann:13 for (p,2p) reactions which includes quantum interference, Eqs. (9) and (11) are purely based on probabilistic concepts and quantum mechanical interference is lost.
Multiple collisions formalisms using eikonal waves have been developed by many authors (see, e.g., PhysRev.103.443 ; PhysRev.184.1745 ; Faeldt1970_DeuteronNucleus ; WALLACE1973190 ; PhysRevLett.32.911 ; RevModPhys.46.279 ; HufnerPRC12.1888 ), although parallel straight-line impact parameters are invoked in the derivations. The model presented here is pure probabilistic, as are other successful models used, e.g., in nucleon-knockout reactions with heavy ion collisions Bertulani:92 ; Hencken:1996 ; Han03 . But it goes beyond by allowing different particle propagation directions after each collision. This idea has been previously applied for (p,2p) reactions in Ref. Aumann:13 although in a different way than it is carried out here because it was intended to describe (p,2p) and use it as a spectroscopic tool in the analysis of reactions involving specific nuclear states. In some sense, when extended to sequential binary collisions the model developed here is similar to intranuclear cascade models. To describe (p,3p) reactions we will use Eq. (9). This requires that, knowing the scattering angle $\Omega$ in the c.m. of the colliding pp pair, a kinematic transformation to the nucleus frame of reference is needed to determine the impact parameter of the leading proton and of the emerging one. To compute the angular functions $W(\Omega)$ we adopt the SAID model for nucleon-nucleon (NN) collisions. It is a partial-wave analysis framework providing NN scattering amplitudes, phase shifts, and observables over a wide range of energies Arndt1987_NN_PWA ; Arndt1994_SAID ; Arndt2007_SAID ; Workman2012_GWU . It is worthwhile mentioning that we did no include a medium modified pp elastic scattering angular distributions. Angular distributions are less critical for inclusive cross sections, and medium effects mainly renormalize attenuation, not angular sampling.
@@
To make calculations tractable, we assume that the reaction is coplanar which simplifies the numerics considerably. Coplanar geometry has been shown to be dominant in (p,2p) reactions Cha77 . This implies a cylindrical symmetry, reducing the number of integrations because the aforementioned impact parameters $b_{1}$ and $b_{2}$ are determined once the scattering angles $\theta_{1}$ and $\theta_{2}$ are known. The remaining integrations are done using a Monte-Carlo sampling over the impact parameter, and positions ${\bf r}_{1}$ and ${\bf r}_{2}$. It is worthwhile mentioning that this procedure is much simpler than that adopted in Ref. Aum13 , which involves a more elaborated calculation with less constraining assumptions. A schematic view of a (p,3p) reaction is shown in Figure 1, where the angles are shown in the nucleus system. For quasi-free elastic scattering and static protons within the nuclei, the angles $\theta_{1L}$ and $\theta_{2L}$ are exactly equal to $90^{\circ}$ because of energy-momentum conservation laws. But they will deviate from this value if one accounts for nuclear Fermi motion or inelastic scattering, not considered in this work. For high-energy protons, the c.m. elastic scattering favors forward scattering which renders the incident proton being scattered in the near forward direction in the nucleus frame with small energy change.
@@ II.1 Numerical Results
Our results for the cross sections are shown in Fig. 2 (open diamonds) compared to the experimental data collected from Ref. PhysRevLett.125.012501 (filled symbols) for (p,2p) and (p,3p) reactions as a function of the nuclear mass. The dashed lines are guides to the eye. Despite the logarithmic scale and far from perfect reproduction of the measured values, the calculations follow the trend of the data with the nuclear mass dependence. In a few cases the calculations differ from the data by an order of magnitude. But in most cases the differences are about a factor 2-5. This level of agreement is encouraging, given the simplicity of the eikonal formalism and the assumption of sequential collisions and coplanar scattering. The coplanarity assumption is known to be well justified for (p,2p) collisions and has been shown to compare well with data Cowley:1995 ; PhysRevC.57.3185 ; Neveling:2002 ; PhysRevC.71.064606 and the interpretation for the good reasonably good agreement is because of a similar amount of absorption along the path of the protons in coplanar and non-coplanar trajectories. In other words, only for off-plane measurements of differential cross sections the non-coplanar feature is of relevance, while for total cross sections, the azimuthal directions of the protons average out. A similar argument may plausibly apply to (p,3p) reactions, although this assumption is less well tested.
@@ III Momentum distribution formalism
A description of the angular and momentum distribution of the recoiled nucleus, such as that published in Ref. Aum13 for (p,2p) reactions, is not appropriate in the case of (p,3p) reactions. We develop here a different approach. The purpose of the calculation is to determine the longitudinal momentum dispersion of projectile fragments produced in sequential proton-proton collisions, for (a) one-proton removal, $(p,2p)$, leaving a residue (Z1,N)(Z-1,N), and (b) two-proton removal, $(p,3p)$, leaving a residue (Z2,N)(Z-2,N) from a tin projectile of mass number $A$ incident on a proton target at 250 MeV, same case as in Ref. PhysRevLett.125.012501 . The observable of interest is the longitudinal (beam-axis) momentum width $\Delta_{\parallel}(A)=\sqrt{\langle P_{z}^{2}\rangle-\langle P_{z}\rangle^{2}},$ where $P_{z}$ is the recoil momentum of the primary fragment. The calculation is strictly limited to sequential collisions with evaporation and ablation explicitly excluded. We only assess the effects imprinted on longitudinal momentum distributions because transverse momentum distributions are known to be more sensitive to additional details of the nucleon-nucleon interactions AumBar21 .
@@
using the functions $K_{\alpha}({\bf b},{\bf r}_{1})$ and $K_{\alpha,\beta}({\bf b},{\bf r}_{1},{\bf r}_{2})$ described in Eqs. (9) and (11). After that a momentum vector ${\bf p}_{i}$ is drawn from the corresponding local Fermi sphere at each point ${\bf r}_{j}$ ($j=1$ or $j=1,2$). The recoil momentum of the primary fragment in (p,2p) reactions is given by

### v2

@@ I Introduction
Fragmentation and knockout reactions, such as (p,2p) (and possibly (p,3p)), at intermediate and high energies could provide a potentially sensitive probe of neutron skins Kox1987 ; Zenihiro2018 . In these reactions, initial- and final-state interactions depend mainly on the total density $\rho=\rho_{p}+\rho_{n}$ and are strongly surface weighted due to absorption and attenuation. Sensitivity to neutron densities is enhanced because $\sigma_{pn}>\sigma_{pp}$ at several hundred MeV to GeV energies, making reaction dynamics increasingly neutron dominated. Glauber-type attenuation, nuclear transparency, and multiple-scattering effects therefore depend primarily on the surface matter distribution rather than on volume-averaged densities. This neutron dominance is most pronounced in the intermediate-energy regime where Glauber and eikonal descriptions are applicable and tends to saturate at asymptotically higher energies. Consequently, incorrect neutron densities can bias extracted spectroscopic factors, nuclear radii, and reaction cross sections by 5–20% or more, primarily through reaction filtering and absorption rather than through bound-state normalization. Glauber and DWIA studies have shown that reaction observables such as $(p,2p)$ cross sections, nuclear transparency, and knockout spectroscopic factors require accurate neutron densities to be interpreted reliably Hufner1985 ; Giusti1999 ; Tostevin1999 ; Hen2017 ; Ogata2015 . While proton-induced knockout reactions probe the nuclear interior, strong attenuation in the entrance and exit channels leads to an effective surface weighting of the reaction amplitude, particularly for neutron-rich nuclei. This is consistent with previous studies showing that $(p,2p)$ reactions retain sensitivity to the interior while exhibiting enhanced sensitivity to surface and density-gradient effects in asymmetric systems aumann:2013:PRC ; Bertulani.104.L061602 .
Momentum distributions are particularly sensitive to the spatial extent of nucleon wave functions, and therefore can reflect surface properties such as neutron skins and halos. In neutron-rich nuclei, valence neutrons may occupy weakly bound orbitals extending into the low-density surface region. The increased spatial localization length leads, via the uncertainty principle, to narrower momentum distributions. As a result, knockout momentum distributions provide a clean probe of density tails and surface diffuseness, complementing information obtained from interaction cross sections or elastic scattering. At higher missing momenta, typically above $p\sim 250$$300\penalty 10000\ \mathrm{MeV}/c$, momentum distributions probe physics beyond the independent-particle picture. In this regime, contributions from short-range and tensor correlations dominate, giving rise to high-momentum tails that are absent in mean-field models. Measurements of these tails, and their isospin dependence in $(p,2p)$ versus $(p,pn)$ reactions, provide direct evidence for correlated nucleon pairs and help explain the systematic quenching of spectroscopic factors observed in knockout experiments Hen2017 ; subedi:2008:SCIENCE .
@@
Ref. aumann:2013:PRC has shown that experimentally observed momentum distributions are intrinsically reaction-dependent quantities, reflecting a nontrivial convolution of nuclear structure and reaction dynamics rather than a simple Fourier transform of the bound-state wave function. Reaction distortions further modify the shape of momentum distributions. Initial- and final-state interactions, recoil effects, and off-shell kinematics can lead to broadening, skewness, or asymmetries, even for nominally symmetric orbitals Giusti1999 ; aumann:2013:PRC ; Ryckebusch2003 . An additional complication arises in the interpretation of high-momentum components. While short-range and tensor correlations in the nuclear wave function generate genuine high-momentum tails, similar features can also emerge from multistep scattering and rescattering processes inherent to the reaction mechanism. Without a consistent reaction model, it is therefore not possible to attribute high-momentum strength uniquely to many-body correlations. This ambiguity is particularly relevant for $(p,2p)$ and $(p,pn)$ reactions at GeV energies, where both correlated dynamics and reaction-induced effects may contribute Hen2017 ; subedi:2008:SCIENCE .
@@ II Cross section formalism
Another problem is that the “straight-line approximation”, implicit in the Glauber method, Eqs. (3-6), is not adequate in the treatment of specific process and their details, such as in (p,2p) and (p,3p) reactions. The physical picture of the most commonly adopted Glauber formalism for proton-nucleus collisions is that the incoming proton travels straight through the nucleus at high energy and at each transverse position $\mathbf{b}$, the probability of interacting with nucleons is described by the product of survival factors $\left[1-\Gamma_{pN}\right]$. Multiple scattering effects arise from the product over all nucleons: the proton can scatter off one, two, or more nucleons during its passage along a straight-line. Although this approximation neglects deflection and quantum interference effects, it has been shown to provide a reasonable leading-order description of inclusive observables at intermediate and high energies ($\gtrsim 200$ MeV/nucleon) Tostevin1999 ; AumannPRL119 ; Teixeira22 .
@@
between two binary collisions happening between positions ${\bf r}_{1}=({\bf b},z_{1})$ and ${\bf r}_{2}=({\bf b},z_{2})$. Between the two positions, the leading particle is assumed to move along a straight-line path with transverse coordinate $\bf b$}. After one collision, the protons are allowed to move in a different direction. As long as their energies are large enough, they can be appropriately described by eikonal wavefunctions. Their wavefunctions may be further influenced by a real mean-field potential $U_{R}$. However, instead of amplitudes, as done in Ref. aumann:2013:PRC , we will work with the probabilistic reaction propagator
@@
The wavefunctions squared, $|\psi_{\alpha}({\bf r})|^{2}$, account for the knockout of a proton from an orbital with quantum numbers defined by $\alpha$. For quasi-free elastic scattering of static protons in the nonrelativistic limit, the sum of the laboratory angles $\theta_{1L}+\theta_{2L}$ equals $90^{\circ}$, with relativistic and Fermi-motion effects leading to deviations from this value. Here we introduce the functions $W(\Omega)$ representing the likelihood for scattering at angles $\Omega$ in the c.m. of the colliding pair. In the case of elastic collisions,
@@
with $d\sigma_{el}/d\Omega$ and $\sigma_{el}$ being the differential and total elastic scattering cross sections, without Coulomb interaction. In summary (see Fig. 1), the impact parameters ${\bf b}_{1}$, and ${\bf b}_{1}^{\prime}$, correspond to vector perpendicular distances between the nuclear center and the velocity vector for the leading proton, and of the proton knocked out from the nucleus, after the first collision. A similar notation, with ${\bf b}_{2}$ and ${\bf b}_{2}^{\prime}$, is used for the second collision. The spectroscopic information about the removed protons are encoded in the wavefunctions $\psi_{i}({\bf r}_{j})$.
This formalism could be easily extended to inelastic collisions such as pion production, by replacing in $W(\Omega)$ the cross sections for pion production, or any other particle produced during the collisions. For (p,2p) collisions, we will use a similar method, with
@@
One can also extend it to more than two binary collisions. In contrast to the method employed in Ref. aumann:2013:PRC for (p,2p) reactions which includes quantum interference, Eqs. (9) and (11) are purely based on probabilistic concepts and quantum mechanical interference is lost.
Multiple collisions formalisms using eikonal waves have been developed by many authors (see, e.g., PhysRev.103.443 ; PhysRev.184.1745 ; Faeldt1970_DeuteronNucleus ; WALLACE1973190 ; PhysRevLett.32.911 ; RevModPhys.46.279 ; HufnerPRC12.1888 ), although parallel straight-line impact parameters are invoked in the derivations. The model presented here is pure probabilistic, as are other successful models used, e.g., in nucleon-knockout reactions with heavy ion collisions Bertulani:92 ; Hencken:1996 ; Han03 . But it goes beyond by allowing different particle propagation directions after each collision. This idea has been previously applied for (p,2p) reactions in Ref. aumann:2013:PRC although in a different way than it is carried out here because it was intended to describe (p,2p) and use it as a spectroscopic tool in the analysis of reactions involving specific nuclear states. In some sense, when extended to sequential binary collisions the model developed here is similar to intranuclear cascade models. To describe (p,3p) reactions we will use Eq. (9). This requires that, knowing the scattering angle $\Omega$ in the c.m. of the colliding pp pair, a kinematic transformation to the nucleus frame of reference is needed to determine the impact parameter of the leading proton and of the emerging one. To compute the angular functions $W(\Omega)$ we adopt the SAID model for nucleon-nucleon (NN) collisions. It is a partial-wave analysis framework providing NN scattering amplitudes, phase shifts, and observables over a wide range of energies Arndt1987_NN_PWA ; Arndt1994_SAID ; Arndt2007_SAID ; Workman2012_GWU . It is worthwhile mentioning that we did no include a medium modified pp elastic scattering angular distributions. Angular distributions are less critical for inclusive cross sections, and medium effects mainly renormalize attenuation, not angular sampling.
@@
We emphasize that the present calculation yields inclusive theoretical cross sections obtained by summing over all occupied proton single-particle orbitals. Removal of deeply bound protons may populate highly excited residual nuclei that subsequently undergo particle evaporation or fragmentation and therefore may not contribute to experimentally identified (A1)(A-1) or (A2)(A-2) residues. Such secondary decay effects are not included in the present framework. While this may affect the absolute normalization of the calculated cross sections, the relative systematics and their dependence on neutron-skin thickness are governed primarily by attenuation and surface geometry and are therefore expected to be robust.
To make calculations tractable, we assume that the reaction is coplanar which simplifies the numerics considerably. Coplanar geometry has been shown to be dominant in (p,2p) reactions Cha77 . This implies a cylindrical symmetry, reducing the number of integrations because the aforementioned impact parameters $b_{1}$ and $b_{2}$ are determined once the scattering angles $\theta_{1}$ and $\theta_{2}$ are known. The remaining integrations are done using a Monte-Carlo sampling over the impact parameter, and positions ${\bf r}_{1}$ and ${\bf r}_{2}$. It is worthwhile mentioning that this procedure is much simpler than that adopted in Ref. aumann:2013:PRC , which involves a more elaborated calculation with less constraining assumptions. A schematic view of a (p,3p) reaction is shown in Figure 1, where the angles are shown in the nucleus system. For quasi-free elastic scattering and static protons within the nuclei, the angles $\theta_{1L}$ and $\theta_{2L}$ are exactly equal to $90^{\circ}$ because of energy-momentum conservation laws. But they will deviate from this value if one accounts for nuclear Fermi motion or inelastic scattering, not considered in this work. Although the proton–proton elastic cross section is symmetric in the center-of-mass frame (after removal of Coulomb effects), transformation to the laboratory (nucleus) frame leads to a forward focusing of the outgoing protons at high energies.
@@ II.1 Numerical Results
Our results for the cross sections are shown in Fig. 2 (open diamonds) compared to the experimental data collected from Ref. PhysRevLett.125.012501 (filled symbols) for (p,2p) and (p,3p) reactions as a function of the nuclear mass. The dashed lines are guides to the eye. Despite the logarithmic scale and far from perfect reproduction of the measured values, the calculations follow the trend of the data with the nuclear mass dependence. In a few cases the calculations differ from the data by an order of magnitude. But in most cases the differences are about a factor 2-5. The pronounced reductions of the $(p,3p)$ cross sections observed around $A=82$ and $A=101$ reflect changes in shell structure and surface density gradients along the isotopic chain. These structural effects modify the effective attenuation and spatial localization of the knockout probability, leading to enhanced suppression of the two-proton removal channel at specific neutron numbers.
The level of agreement of our calculations with experimental data is encouraging, given the simplicity of the eikonal formalism and the assumption of sequential collisions and coplanar scattering. The coplanarity assumption is known to be well justified for (p,2p) collisions and has been shown to compare well with data Cowley:1995 ; PhysRevC.57.3185 ; Neveling:2002 ; PhysRevC.71.064606 and the interpretation for the good reasonably good agreement is because of a similar amount of absorption along the path of the protons in coplanar and non-coplanar trajectories. In other words, only for off-plane measurements of differential cross sections the non-coplanar feature is of relevance, while for total cross sections, the azimuthal directions of the protons average out. A similar argument may plausibly apply to (p,3p) reactions, although this assumption is less well tested.
@@
Semiclassical and eikonal descriptions of multistep nucleon-induced reactions have a long history. Early work by Kawai and Weidenmüller KawaiWeidem developed a semiclassical treatment of two-step nucleon-nucleus processes that is closely related in spirit to the probabilistic multistep approach adopted here. More recently, an eikonal description of (p,3p) reactions was presented by Gómez-Ramos in Ref. Ramos2024 . While both approaches employ eikonal dynamics, the present framework emphasizes inclusive observables and density-driven attenuation effects rather than exclusive final-state amplitudes. Benchmarks for (p,pN) reactions have also been studied in Refs. RDM2020 ; YoshRam2018 .
@@ III Momentum distribution formalism
A description of the angular and momentum distribution of the recoiled nucleus, such as that published in Ref. aumann:2013:PRC for (p,2p) reactions, is not appropriate in the case of (p,3p) reactions. We develop here a different approach. The purpose of the calculation is to determine the longitudinal momentum dispersion of projectile fragments produced in sequential proton-proton collisions, for (a) one-proton removal, $(p,2p)$, leaving a residue (Z1,N)(Z-1,N), and (b) two-proton removal, $(p,3p)$, leaving a residue (Z2,N)(Z-2,N) from a tin projectile of mass number $A$ incident on a proton target at 250 MeV, same case as in Ref. PhysRevLett.125.012501 . The observable of interest is the longitudinal (beam-axis) momentum width $\Delta_{\parallel}(A)=\sqrt{\langle P_{z}^{2}\rangle-\langle P_{z}\rangle^{2}},$ where $P_{z}$ is the recoil momentum of the primary fragment. The calculation is strictly limited to sequential collisions with evaporation and ablation explicitly excluded. We only assess the effects imprinted on longitudinal momentum distributions because transverse momentum distributions are known to be more sensitive to additional details of the nucleon-nucleon interactions AumBar21 .
@@
using the functions $K_{\alpha}({\bf b},{\bf r}_{1})$ and $K_{\alpha,\beta}({\bf b},{\bf r}_{1},{\bf r}_{2})$ described in Eqs. (9) and (11). After that a momentum vector ${\bf p}_{i}$ is drawn from the corresponding local Fermi sphere at each point ${\bf r}_{j}$ ($j=1$ or $j=1,2$). Here $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ denote the spatial coordinates of the first and second knocked-out protons, respectively, measured from the center of the projectile nucleus, and $\mathbf{r}\equiv(\mathbf{r}_{1},\mathbf{r}_{2})$ is used as a shorthand notation in the (p,3p) case. The recoil momentum of the primary fragment in (p,2p) reactions is given by